How many terms are in the sequence: 640,320,160,.......,5?
Found 2 solutions by macston, dkppathak:
Answer by macston(5194)

640,320,160,80,40,20,10,5
There are 8 terms in the series

Answer by dkppathak(439)

How many terms are in the sequence: 640,320,160,.......,5?
it is gp whose a=640 r=1/2=2^-1 n=? last terms is=5
last term =ar^n-1
5=640x(1/2)^n-1
5/640 =2^-n+1
1/128=2^-n+1
2^-7=2^-n+1
-7=-n+1
-7-1=-n
n=8